# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ 3.2 Meta-Learning
In this section, we generalize the in-distribution bilevel optimization framework to tackle the out-of-distribution (OOD) preference learning. To this end, we assume that there exists a meta-distribution $\mathcal{D}_{\mathcal{T}}$ over the set of distributions $\mathcal{T}$, e.g., $\mathcal{T}:=\{{\tt book},{\tt dvd},{\tt film},\ldots\}$. Previous study on the distribution-level generalization refers to meta-learning (Baxter, 2000; Vilalta & Drissi, 2002; Finn et al., 2017).
@@
In order to learn a generalized reward modeling for fine-tuning unseen distribution in the test phase, we utilize a meta-learning framework for reward optimization. To this end, the outer objective described in the Section 3.1 should be extent to computing the preference loss over multiple distributions, while the inner objective can be unchangeable as a fine-tuning objective for each training distribution. In particular, given a distribution $T\in\mathcal{T}$, there exists a distribution-specific fine-tuning sampling distribution $\mathcal{D}^{T}$. For preference learning, each training distribution has a preference sampling distribution $z=(x,y,y^{\prime})\sim\mathcal{D}^{T}_{\rm PL}$. The optimal variable in inner objective should be the distribution-specific policy ${\theta^{T}_{*}}$. A key point is that the reward model ${\phi}$ is shared across all distributons. With this notation the inner and outer objectives can be represented as follows.
$\min_{\phi}{\bf L}(\phi)=\min_{\phi}\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}_{z\sim\mathcal{D}^{T}_{\rm PL}}\left[\ell_{\rm PL}\left(\phi,\theta^{T}_{*},z\right)\right]$ (11)
${\rm s.t.}\ \theta^{T}_{*}\in\mathop{\arg\min}_{{\theta^{T}}}\mathbb{E}_{z\sim\mathcal{D}^{T}}\left[\ell_{\rm PL}({\phi},{\theta^{T}},z)\right],$ (12)
@@
where the objective function $\ell_{\rm PL}(\phi,{\theta^{T}_{*}},z)$, $z\sim\mathcal{D}^{T}_{\rm PL}$ represents the preference learning loss, which has a same formulation as the in-distribution version of Eq. (7), and $\ell_{\rm FT}(\phi,\theta^{T},z)$, $z\sim\mathcal{D}^{T}$ represents the fine-tuning loss, which has the same formulation as the in-distribution version of Eq. (10).
@@ 3.3 Gradient-Based Algorithm
Input: Meta-training data $\{\mathcal{D}^{T}_{\rm PL},\mathcal{D}^{T}\}_{T\in\mathcal{T}_{tr}}$, reward model $\phi$, policy $\theta$
Output: Optimal reward model $\phi_{*}$ for meta-test in Algorithm 2
for $k=0,1,2,\ldots,K$ do
       Randomly draw a sample distribution $T\in\mathcal{T}_{tr}$ from the meta-distribution $\mathcal{D}_{\mathcal{T}}$
       Parameter initialization $\theta\leftarrow\theta_{0}$
       for $t=0,1,2,\ldots,D$ do
             Randomly draw a sample $z_{k,t}$ from $\mathcal{D}^{T}$
@@
      Randomly draw a sample $z_{k}$ from $\mathcal{D}^{T}_{\rm PL}$
@@
Input: Meta-test data $\mathcal{D}^{te}$; optimized reward function $\phi_{*}$ from Algorithm 1
@@
Parameter initialize $\theta\leftarrow\theta_{0}$
@@
       Randomly draw a sample $z$ from $\mathcal{D}^{te}$
@@
Following the previous work on gradient-based bilevel programming (Finn et al., 2017; Franceschi et al., 2018; Ji et al., 2021), we use a gradient-based approach to tackle the bilevel optimization objective in Eq. (11) and Eq. (12). The whole training process consists of a meta-training procedure and a meta-test procedure.
Meta-training.
As shown in Algorithm 1, for each iteration $k\in{1,\ldots,K}$ in the outer-loop optimization, the inner-loop process runs $D$ steps of stochastic gradient decent (SGD) to gain an optimal solution of ${\theta}_{*}$ with $\theta_{k,D}$. Then, the outer-loop process uses SGD optimization w.r.t. $\phi_{k}$ via computes the gradient $\mathbb{E}\left[\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}\right]$ as an approximation of the hyper-gradient $\mathbb{E}\left[\frac{\partial\ell_{\rm PL}(\phi_{k},{\theta}_{*})}{\partial\phi_{k}}\right]$. Specifically, we compute the explicit form of the $\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}$ in the following proposition.
Proposition 1.
For any outer-loop step $k\in\{0,1,2,\ldots,K-1\}$, with the outer-loop input $z_{k}$ and the inner-loop inputs $\{z_{k,t}\}_{t=0}^{D-1}$, the gradient w.r.t. $\phi_{k}$ $\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}$ takes the analytical form,
$\begin{split}\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial{\phi}_{k}}=&\frac{\alpha}{\beta}\left(\nabla_{\theta}\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\nabla_{\theta}\log\pi_{\theta_{k,D}}(y^{\prime}_{k}|x_{k})\right)\\ &\cdot\sum_{t=0}^{D-1}\bigg{\{}\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)\nabla_{\theta}\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})\\ &\cdot\prod_{j=t+1}^{D-1}\left(I+\alpha\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)\nabla_{\theta}^{2}\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})\right)\bigg{\}}\end{split}$ (13)
Proof.
The major proof steps consists of the chain rule of partial differentiation, and the detailed derivation is available in Appendix B. ∎
Finally, the meta-training procedure outputs an optimized reward function parameterized by $\phi_{*}$.
Meta-test.
As shown in Algorithm 2, the meta-test procedure takes the optimal reward model $\phi_{*}$ from meta-training to learn a policy for a given test distribution. In particular, the meta-test process computes policy gradients based on the optimal reward model $\phi_{*}$ and uses SGD to optimize the distribution-specific policy with a sufficient number of iterations for convergence. The meta-test procedure can also be viewed as test-phase adaptation.
@@ 3.4 Convergence Analysis
Following previous work on convergence analysis of SGD (Drori & Shamir, 2020; Ji et al., 2021), we study the convergence rate of the proposed SGD-based bilevel programming to the stationary point. In particular, we aim to show how the proposed algorithm minimizes $\mathbb{E}[\|\nabla_{\phi}{\bf L}(\overline{\phi})\|^{2}]$, where ${\bf L}(\phi)$ is defined in Eq. (11) and $\overline{\phi}$ denotes the output of the algorithm.
Following previous work on the convergence rate of gradient-based bilevel optimization (Ji et al., 2021), we focus on the following types of basic loss functions w.r.t. the policy for preference learning.
Assumption 1.
@@
Besides, we take the following standard Lipschitz assumptions on the basic MLE loss functions of policy $-\log\pi_{\theta}$ and the reward function $r_{\phi}$, which are largely followed those have been widely adopted in bilevel optimization (Ghadimi & Wang, 2018; Ji et al., 2021).
Assumption 2.
The loss function $-\log\pi_{\theta}$ and reward function $r_{\phi}$ satisfy
@@
Besides, we also consider the following reasonable assumptions on the boundnesses of reward function and predicted probability by the policy.
Assumption 3.
The function $-\log\pi_{\theta}(y|x)$ and $\exp(r_{\phi}(x,y))$ are bounded, i.e., for any $(x,y)$, there exist $b,B,C>0$,
$b\leq\exp(r_{\phi}(x,y))\leq B$ (19)
$|\log\pi_{\theta}(y|x)|\leq C$ (20)
As typically adopted in the analysis of stochastic optimization by Ji et al. (2021), we make the following assumption of bounded variances for the loss function.
Assumption 4.
Gradients $\nabla_{\theta}\ell_{\rm FT}({\phi},{\theta};z)$, $\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(r_{\phi},\pi_{\theta};z)$, $\nabla^{2}_{\theta}\ell_{\rm FT}(r_{\phi},\pi_{\theta};z)$ and $\frac{\partial\ell_{\rm PL}(r_{\phi},\pi_{{\theta}};z)}{\partial\phi}$ have bounded variances,
@@
$\mathbb{E}_{z}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi},{\theta})}{\partial\phi}-\frac{\partial\ell_{\rm PL}({\phi},{\theta};z)}{\partial\phi}\right\|^{2}\right]\leq$ $\sigma^{2}_{3}$ (24)
Then, we present the main convergence result in this work based on the above assumptions.
Theorem 1.
Suppose Assumption 1, 2, 3 and 4 hold, choose the iteration number $D$ in the inner-loop to satisfy $D\geq-\frac{\log\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)}{\log\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)}$, and assume that the reward controlling factor satisfies $\beta\leq\min\Big{\{}\log\frac{3\alpha B^{2}\left(L^{2}+\frac{M^{2}\rho^{2}}{L^{2}}\right)}{2b\mu\left(3-\frac{BL}{b\mu+BL}\right)},\log\left(\alpha b\mu/1-\sqrt{\frac{1-\kappa}{3}}\right)\Big{\}}$, for some coefficient $\kappa>0$, then we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{K}+\frac{1}{\beta^{2}}\right)+\tau,$ (25)
where $\tau$ denotes a $K$, $\beta$-unrelated term, see Appendix 1 for a detailed formulation.
Proof.
See Appendix D for detailed derivations. ∎
Remark 1.
Theorem 1 shows that the proposed bilevel optimization algorithm converges w.r.t. the number of outer-loop iteration $\mathcal{O}\left(\frac{1}{K}\right)$ and the controlling factor of reward function $\mathcal{O}\left(\frac{1}{\beta^{2}}\right)$. This suggest that (i) more iterations of SGD and (ii) larger reward controlling factor makes the algorithm converge better.
4 Experiments
We conduct experiments on two text generation tasks and use three types of metrics to evaluate the performance of the proposed approach.
@@ 4.1 Experimental Setup
For controlled sentiment generation, $x$ represents a prefix of an amazon review from the amazon review dataset, which contains four domains (Blitzer et al., 2007). The policy generates the remaining review context $y$ with a positive sentiment. It is noted that the original dataset is intended for text classification and contains only a single review for each prefix. To construct a pairwise dataset for preference learning, we follow the approach of previous work, RLAIF (Lee et al., 2023), using GPT-4 (Achiam et al., 2023) to generate pairwise reviews for each prefix. Detailed prompts for generating the preference data are provided in Appendix E.1. Following Li et al. (2017), we adopt a leave-one-domain-out protocol for out-of-distribution (OOD) evaluation in our experiments. Specifically, each experiment on amazon reviews is trained on three of the four distributions and evaluated on the remaining distribution. For instance, the training set includes dvd, electronics, and kitchen, while the testing set consists of book reviews. For the knowledge answer generation task, we utilize the Stanford Human Preferences Dataset (SHP) (Ethayarajh et al., 2022), which comprises 18 domains of knowledge question-answering (QA) with preference labels. Accordingly, $x$ is a Reddit question, the policy generates a proper response $y$ to the question. To evaluate the OOD performance, we employ a cross-validation evaluation method on the SHP dataset. Specifically, we randomly divide the 18 domains of the SHP dataset into six splits, with each split containing three domains. In each experiment, one split is used for testing, while the remaining splits are used for training. For instance, the training set includes ca,cu,do,ne,hi,hr,phi,sci,scif,so,ve,ch,ex,le, and the testing set consists of ac,an,ba.
Metrics.
@@
We use three metrics for evaluation. To assess the effectiveness of policy learning in aligning with human preferences, we introduce a metric called PL accuracy $\mathcal{A}_{\rm PL}$, which quantifies the proportion of correctly predicted comparisons using the normalized language probabilities generated by the learned policy across all pairwise comparisons in the preference test set.
@@
$\mathcal{A}_{\rm PL}(\theta;\mathcal{D}_{\rm PL})=\frac{1}{|\mathcal{D}_{\rm PL}|}\sum_{(y,y^{\prime},x:y\succ y^{\prime}|x)\in\mathcal{D}_{\rm PL}}\mathbb{I}\left[\frac{\log\pi_{\theta}(y|x)}{|y|}>\frac{\log\pi_{\theta}(y^{\prime}|x)}{|y^{\prime}|}\right],$ (26)
[…]
Optimization Protocol.
@@
For model selection, we select the best reward model $\phi_{*}$ learned in the meta-training phase based on the minimum PL accuracy on the evaluation sets, and choose the best policy model $\theta_{*}$ learned in the meta-test phase based on the minimum training loss. The hyperparameters used in our experiments largely follow previous works (Ramamurthy et al., 2023; Ouyang et al., 2022). Additionally, we select hyperparameters related to bilevel programming from a preset range based on development experiments. For example, the inner-loop learning rate $\alpha$ is chosen from ${5e^{-6},1e^{-5},3e^{-5},5e^{-5}}$, the reward controlling factor $\beta$ is selected from ${0.1,1.0,2.0,4.0,8.0}$, and the inner-loop iteration number $D$ is selected from ${50,100,150,200}$.
Approaches.
@@
Our methods are built upon the Base Model Flan-T5 XXL Chung et al. (2022), an enhanced version of T5 (Raffel et al., 2020) fine-tuned on a mixture of tasks, comprising more than 11B parameters. To expedite the training procedure, we adopt the parameter-efficient fine-tuning method LoRA (Hu et al., 2021). Specifically, the trainable parameters of a reward model consist of a LoRA adapter and a value head, while the trainable parameters of a policy consist of a LoRA adapter and a LM head. In addition to the proposed method, we also consider several strong baselines for out-of-distribution (OOD) preference learning. For Sft, we utilize supervised data from both the training and test distributions for policy fine-tuning. The other approaches use SFT parameters for model initialization. For a fair comparison with PPO (Schulman et al., 2017), we use preference data from all training distributions to train the reward model with a multi-task learning approach. We also consider two RLHF approaches: Ppo (tr) uses supervised data from only training distributions for policy fine-tuning, while Ppo (tr+te) uses supervised data from both training and test distributions for policy fine-tuning. For Dpo (Rafailov et al., 2024), we use preference data from the training distributions for policy training. Additionally, we also consider a variation of our approach named Hpl, which directly optimizes the preference learning objective $\mathcal{L}_{\rm PL}$ using preference data from both the training and test distributions.

### v2

@@ 3.2 Meta-Learning for Out-of-Distribution Preference Learning
We generalize the ID bilevel optimization framework to tackle OOD preference learning. To this end, we assume that there exists a meta-distribution $\mathcal{D}_{\mathcal{T}}$ over the set of tasks $\mathcal{T}$. Previous studies on the distribution generalization refer to meta-learning [5, 16].
@@
To learn a generalized reward modeling for RL fine-tuning on unseen test distributions, we utilize a meta-learning framework, where the outer objective described in the §3.1 should be the extent to compute the preference loss across tasks, and the inner objective can be unchangeable for each distribution. In particular, given a task $T\in\mathcal{T}$, there exists a task-specific distribution $\mathcal{D}^{T}_{xy}$. For preference learning, each training distribution has a preference distribution $\nu=(x,y,y^{\prime})\sim\mathcal{D}^{T}_{\rm PL}$. The optimal variable in inner objective should be the task-specific policy ${\theta^{T}_{*}}$. A key point is that the RM $r_{\phi}$ is shared across all task distributions. With this notation, the inner and outer objectives can be represented as follows,
$\min_{\phi}{\bf L}(\phi)=\min_{\phi}\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}_{\nu\sim\mathcal{D}^{T}_{\rm PL}}\left[\ell_{\rm PL}\left(\phi,\theta^{T}_{*},\nu\right)\right]$ (11)
${\rm s.t.}\ \theta^{T}_{*}\in\mathop{\arg\min}_{{\theta^{T}}}\mathbb{E}_{z\sim\mathcal{D}^{T}_{xy}}\left[\ell_{\rm FT}({\phi},{\theta^{T}},z)\right],\vspace{-0.5cm}$ (12)
@@
where the objective $\ell_{\rm PL}(\phi,{\theta^{T}_{*}},\nu)$, $\nu\sim\mathcal{D}^{T}_{\rm PL}$ represents the preference learning loss, which has a same formulation as the ID version of Eq. (7), $\ell_{\rm FT}(\phi,\theta^{T},z)$, $z\sim\mathcal{D}^{T}_{xy}$ represents the fine-tuning loss, which has the same formulation as the ID version of Eq. (10).
@@ 3.3 Gradient-Based Algorithm
Input: Meta-training data $\{\mathcal{D}^{T}_{\rm PL},\mathcal{D}^{T}_{xy}\}_{T\in\mathcal{T}_{tr}}$, RM $\phi$, policy $\theta$
Output: Optimal RM $\phi_{*}$ for meta-test in Alg. 2
for $k=0,1,2,\ldots,K-1$ do
       Randomly draw a sample task $T\in\mathcal{T}_{tr}$ from the meta-distribution $\mathcal{D}_{\mathcal{T}}$
       Initialize $\theta$ with SFT parameters on $\mathcal{D}^{T}_{xy}$
       for $t=0,1,2,\ldots,D-1$ do
             Randomly draw a sample $z_{k,t}$ from $\mathcal{D}^{T}_{xy}$
@@
      Randomly draw a sample $\nu_{k}$ from $\mathcal{D}^{T}_{\rm PL}$
@@
Input: Meta-test data $\mathcal{D}^{te}_{xy}$; optimized reward function $\phi_{*}$ from Alg. 1
@@
Initialize $\theta$ with SFT parameters on $\mathcal{D}^{te}_{xy}$
@@
       Randomly draw a sample $z$ from $\mathcal{D}^{te}_{xy}$
@@
Following the previous work on gradient-based bilevel programming [16, 19], we use a gradient-based approach to tackle the bilevel optimization objective in Eq. (11) and Eq. (12). The whole training process consists of meta-training and meta-test procedures, as shown in Fig. 2.
Meta-training. As illustrated in Alg. 1, for each outer-loop iteration $k\in\{0,1,\ldots,K-1\}$, the inner-loop process runs $D$ steps of stochastic gradient decent (SGD) to obtain $\theta_{k,D}$. Then, the outer-loop process uses SGD optimization w.r.t. $\phi_{k}$ via computes the gradient $\mathbb{E}\left[\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}\right]$. Specifically, we compute the explicit form of the $\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}$ in the following proposition. After $K$ steps of meta-training, Alg. 1 outputs optimized parameters of the reward function $\phi_{*}$.
Proposition 1
For any outer-loop step $k\in\{0,1,2,\ldots,K-1\}$, with the outer-loop input $z_{k}$ and the inner-loop inputs $\{z_{k,t}\}_{t=0}^{D-1}$, the gradient $\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial\phi_{k}}$ takes the analytical form,
$\footnotesize\vspace{-0.5cm}\begin{split}\frac{\partial\ell_{\rm PL}(\phi_{k},\theta_{k,D})}{\partial{\phi}_{k}}=-\frac{\alpha}{\beta}\nabla_{{\theta}}A_{k}\left(1-\sigma\left(A_{k}\right)\right)\sum_{t=0}^{D-1}\nabla_{\phi}H_{k,t}\nabla_{{\theta}}^{\top}F_{k,t}R_{k,t}\prod_{j=t+1}^{D-1}\left(I+\alpha\nabla_{{\theta}}^{2}F_{k,t}R_{k,t}\right),\end{split}\vspace{-0.5cm}$
where $A_{k}:=\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\log\pi_{{\theta}_{k,D}}(y^{\prime}_{k}|x_{k})$, $R_{k,t}:=\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)$, $F_{k,t}:=\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})$ and $H_{k,t}:=r_{\phi_{k}}(x_{k,t},y_{k,t})$.
Proof
The major proof steps consist of the chain rule of partial differentiation, and the detailed derivation is available in Appendix 0.B.
Meta-test. As illustrated in Alg. 2, the meta-test procedure takes the optimal RM $\phi_{*}$ by meta-training to learn a policy for a given test distribution. In particular, the meta-test process computes policy gradients based on the optimal RM $\phi_{*}$ and uses SGD to optimize the task-specific policy with a sufficient number of iterations for convergence. The meta-test procedure can also be viewed as test-phase adaptation.
@@ 3.4 Convergence Analysis
Following previous work on convergence analysis of SGD [13, 19], we study the convergence rate of the proposed bilevel programming to the stationary point. In particular, we aim to show how the proposed algorithm minimizes $\mathbb{E}[\|\nabla_{\phi}{\bf L}(\overline{\phi})\|^{2}]$, where ${\bf L}(\phi)$ is defined in Eq. (11) and $\overline{\phi}$ denotes the output of algorithm. Following previous work on the convergence rate of gradient-based bilevel optimization [19], we focus on the following types of basic loss functions for PL.
Assumption 1
@@
Besides, we take the following standard Lipschitz assumptions on the basic MLE loss functions of policy $-\log\pi_{\theta}$ and the reward function $r_{\phi}$, which are largely followed by those that have been widely adopted in bilevel optimization [17, 19].
Assumption 2
The loss function $-\log\pi_{\theta}$ and reward function $r_{\phi}$ satisfy (1) $-\log\pi_{\theta}(y|x)$ is $M$-Lipschitz. (2)$-\nabla\log\pi_{\theta}(y|x)$ is $L$-Lipschitz. (3) $-\nabla^{2}\log\pi_{\theta}(y|x)$ is $\rho$-Lipschitz. (4) $r_{\phi}$ is $T$-Lipschitz. (5) $\nabla r_{\phi}(x,y)$ is $P$-Lipschitz. More specifically,
@@
We also consider the following reasonable assumptions on the boundness of reward function and policy prediction probability.
Assumption 3
The function $-\log\pi_{\theta}(y|x)$ and $\exp(r_{\phi}(x,y))$ are bounded, i.e., for any $(x,y)$, there exist $b,B,C>0$,$b\leq\exp(r_{\phi}(x,y))\leq B,|\log\pi_{\theta}(y|x)|\leq C$.
As typically adopted in the analysis of stochastic optimization by Ji et al. [19], we make the following assumptions of bounded variances for the loss functions.
Assumption 4
Gradients $\nabla_{\theta}\ell_{\rm FT}({\phi},{\theta};z)$, $\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(r_{\phi},\pi_{\theta};z)$, $\nabla^{2}_{\theta}\ell_{\rm FT}(r_{\phi},\pi_{\theta};z)$ and $\frac{\partial\ell_{\rm PL}(r_{\phi},\pi_{{\theta}};z)}{\partial\phi}$ have bounded variances of $\sigma$,$\sigma^{2}_{1}$,$\sigma^{2}_{2}$,$\sigma^{2}_{3}$, respectively. More specifically,
@@
$\mathbb{E}_{\nu}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi},{\theta})}{\partial\phi}-\frac{\partial\ell_{\rm PL}({\phi},{\theta};\nu)}{\partial\phi}\right\|^{2}\right]\leq$ $\sigma^{2}_{3}$ (21)
Then, we present the main convergence result based on the above assumptions.
Theorem 3.1
Suppose Assumption 1, 2, 3 and 4 hold, choose the iteration number $D$ in the inner-loop and the reward controlling factor $\beta$ satisfy the conditions: (i) $D\geq-\frac{\log\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)}{\log\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)}$, and (ii) $\beta\leq\min\Big{\{}\log\frac{4\alpha B^{2}(L^{4}+M^{2}\rho^{2})}{b\mu(2L+M^{2})^{2}},\log\left(\alpha b\mu/1-\sqrt{\frac{1-\kappa}{3}}\right)\Big{\}}$, for some coefficient $\kappa>0$, then we have
$\vspace{-0.5cm}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{K}+\frac{1}{\beta^{2}}\right)+\tau,\vspace{-0.5cm}$ (22)
where $\tau$ denotes a $K$, $\beta$-unrelated term, which has the form of $\tau=3\sigma_{3}^{3}+\frac{144M^{2}\alpha^{2}\sigma_{1}^{2}}{\kappa}$.
Proof
See Appendix 0.D for detailed derivations.
Remark 1
Theorem 3.1 shows that the proposed bilevel optimization algorithm converges w.r.t. the number of outer-loop iteration $\mathcal{O}\left(\frac{1}{K}\right)$ and the controlling factor of reward function $\mathcal{O}\left(\frac{1}{\beta^{2}}\right)$. This suggests that (a) more iterations of SGD and (b) a larger reward controlling factor under the condition (ii) makes the algorithm converge better.
@@ 3.5 Experimental Setup
Tasks. For the controlled sentiment generation (CSG) task, $x$ represents a prefix of a review from the Amazon Review dataset, which contains four domains [6]. The policy generates the remaining review context $y$ with a positive sentiment. We use GPT-4 [1] to generate pairwise reviews for each prefix. Following Li et al. [21], we adopt a leave-one-out protocol for OOD evaluation. Specifically, each experiment on Amazon Review is trained on three of the four distributions and evaluated on the remaining distribution. For the knowledge answer generation (KAG) task, we utilize the Stanford Human Preferences Dataset (SHP) [14], which comprises 18 domains of knowledge question-answering (QA) with preference labels. Accordingly, $x$ is a Reddit question, the policy generates a proper response $y$ to the question. We employ a cross-validation evaluation method, i.e., randomly dividing the 18 domains of the SHP dataset into six splits, one split is used for testing, while the remaining splits are used for training in each experiment.
@@
Metrics. We use three metrics for evaluation. To assess the effectiveness of policy learning in aligning with human preferences, we introduce a metric called PL accuracy $\mathcal{A}_{\rm PL}$, which quantifies the proportion of correctly predicted comparisons by the learned policy in the preference test set.
@@
$\vspace{-0.5cm}\mathcal{A}_{\rm PL}(\theta;\mathcal{D}_{\rm PL})=\frac{1}{|\mathcal{D}_{\rm PL}|}\sum_{(y,y^{\prime},x:y\succ y^{\prime}|x)\in\mathcal{D}_{\rm PL}}\mathbb{I}\left[\frac{\log\pi_{\theta}(y|x)}{|y|}>\frac{\log\pi_{\theta}(y^{\prime}|x)}{|y^{\prime}|}\right],\vspace{-0.5cm}$ (23)
where the normalization terms $1/|y|$ and $1/|y^{\prime}|$ are incorporated to mitigate the influence of generation length on the computation of language probability. In the CSG task, we assess the quality of generation for preference alignment using a learned RM. Specifically, we fine-tune a RoBERTaLARGE model [23] on the Amazon Review dataset to serve as the ground truth RM for scoring the generated reviews. In the KAG task, as the ground truth reward function is unknown, we evaluate approaches based on their win rate against the preferred responses in the test set. We use GPT-4 judgement to assess response correctness and usefulness.
@@
Optimization Protocol. For model selection, we select the best RM $\phi_{*}$ learned in the meta-training phase based on the minimum PL accuracy on the evaluation sets, and choose the best policy model $\theta_{*}$ learned in the meta-test phase based on the minimum training loss. The hyperparameters used in our experiments largely follow previous works [31, 26]. Additionally, we select hyperparameters related to bilevel programming from a preset range based on development experiments. For example, the inner-loop learning rate $\alpha$ is chosen from ${5e^{-6},1e^{-5},3e^{-5},5e^{-5}}$, the reward controlling factor $\beta$ is selected from ${0.1,1.0,2.0,4.0,8.0}$, and the inner-loop iteration number $D$ is selected from ${50,100,150,200}$.
@@
Approaches. Our methods are built upon the Base Model Flan-T5 XXL [11], comprising more than 11B parameters. To expedite the training procedure, we adopt the parameter-efficient fine-tuning method LoRA [18]. Specifically, the trainable parameters of an RM consist of a LoRA adapter and a value head, while the trainable parameters of a policy consist of a LoRA adapter and a LM head. We also consider several strong baselines of OOD PL for comparison. For SFT, we utilize supervised data from both the training and test distributions for policy fine-tuning. The other approaches use SFT parameters for model initialization. For a fair comparison with PPO [33], we use preference data from all training distributions to train the RM with a multi-task learning approach. We also consider two RLHF approaches: PPO (tr) and PPO (te) use supervised data from the training distributions and test distribution for policy fine-tuning, respectively. For DPO [30], we use preference data from the training distributions for policy training. Additionally, we also consider a variation of our approach named HPL, which directly optimizes the PL objective $\mathcal{L}_{\rm PL}$ using preference data from the training distributions.